OpenStudy (anonymous):

You are camping at the Grand Canyon. When you pitch your tent at 1:00 P. M. the temperature is 81° F. When you go to bed at 11 P.M. the temperature is 61° F. What is the average rate of change in the temperature?

OpenStudy (asadkarim7):

no of hours = 11-1=10 change in temp = -20 so = -20/10 -2 deg F per hour